Circular No. 8046

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003I IN IC 2481 Further to IAUC 8043, T. Puckett and A. Langoussis report the discov-

ery of an apparent supernova (mag 17.6) on an unfiltered CCD frame taken with the 0.60-m automated patrol telescope on Jan. 9.31 UT (confirmed with CCD frames taken Jan. 10.29 by A. Sehgal, Woodinville, WA, 0.35-m reflector). The new object is located at α = 9^{h}27^{m}29^{s}.48, δ = +3^{o}55′45′′.6 (equinox 2000.0), which is 10′′.3 east and 1′′.3 south of the center of IC 2481. SN 2003I is not present on images taken by Puckett on 2001 Mar. 15 and 26 (limiting mag ∼ 20.5) or on Palomar Sky Survey plates taken on 1994 Dec. 3 (limiting mag ∼ 21.1), 1991 Feb. 6, and 1949 Sept. 30.

SUPERNOVA 2003H IN NGC 2207 On IAUC 8045, the offset was given erroneously for IC 2163, the eastern

member of the galaxy pair; SN 2003H is located 51′′.2 east and 2′′.1 south of the nucleus of NGC 2207, the western member of the galaxy pair.

COMET C/2002 X5 (KUDO-FUJIKAWA) Ephemeris extension to IAUC 8033 (from elements on MPC 47293):

2003 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Jan. 11 19^{h}11.^{m}19 +14^{o}06.^{′}9 1.011 0.621 36.^{o}3 69.^{o}5 5.5
16 19 36.97 + 6 12.8 1.055 0.489 27.4 68.1 4.8
21 20 00.24 − 2 21.5 1.114 0.350 17.9 59.6 3.8
26 20 23.05 −12 29.5 1.169 0.223  6.7 31.1 2.5
31 20 51.45 −24 22.7 1.147 0.205  6.8 34.7 2.1
Feb. 5 21 31.02 −33 34.7 1.045 0.322 17.9 70.5 3.4
10 22 23.29 −39 38.7 0.952 0.461 27.4 80.5 4.4
15 23 28.38 −42 54.6 0.890 0.595 36.5 80.8 5.1
20 0 40.12 −42 58.0 0.863 0.721 45.2 76.6 5.6
25 1 47.40 −40 00.3 0.873 0.841 53.2 70.5 6.1
Mar. 2 2 42.74 −35 09.6 0.916 0.955 59.9 63.9 6.7
7 3 25.47 −29 43.2 0.988 1.064 65.0 57.7 7.2
12 3 58.23 −24 31.0 1.082 1.169 68.4 52.2 7.7

Visual m_{1} estimates: 2002 Dec. 16.58 UT, 7.8 (S. J. O’Meara, Volcano, HI, 0.10-m refractor); 24.69, 7.4 (B. H. Granslo, Malm, Norway, 12×50 binoculars); 2003 Jan. 3.25, 6.4 (J. J. Gonzalez, Asturias, Spain, 7×50 binoculars); 7.57, 6.3 (C. S. Morris, Fillmore, CA, 10×50 binoculars).

2003 January 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green